OpenStudy (butterflyhope):

Check *will medal*: 7/12 + 5/9 - 7/8. The LCM of 12, 9, 8 is 72. So: 7/12 = 12 x 6 = 72 and 7 x 6 = 42. So 42/72. 5/9 = 9 x 8 = 72 and 5 x 8 = 40. So 40/72. 7/8 = 8 x 9 = 72 and 7 x 9 = 63. So 63/72. 42/72 + 40/72 - 63/72 = 19/72. Thanks c:

OpenStudy (igreen):

Yeah, you have to convert them all into fractions with a denominator of 72. \(\dfrac{7}{12} + \dfrac{5}{9} - \dfrac{7}{8} \rightarrow \dfrac{42}{72} + \dfrac{40}{72} - \dfrac{63}{72}\)

OpenStudy (butterflyhope):

Thanks c:

OpenStudy (igreen):

\(\dfrac{7}{12} + \dfrac{5}{9} - \dfrac{7}{8} \rightarrow \dfrac{42}{72} + \dfrac{40}{72} - \dfrac{63}{72} \rightarrow \dfrac{82}{72} - \dfrac{63}{72} \rightarrow \dfrac{19}{72}\)

OpenStudy (igreen):

19/72 is your final answer..

OpenStudy (igreen):

I don't exactly know what the question is asking..